can you please help me solve this equation...
10r-9s=18
6s+2r=1
its either a substitution problem or you add or subtract..
i have tried and tried to get this answer but i can't seem to get it..
thank you so much for helping!

Answer by user_dude2008(1862)

6s+2r=1

2r=1-6s

10r=5-30s

---------

10r-9s=18

5-30s-9s=18

5-39s=18

-39s=13

s=-1/3 <--------Answer # 1


10r=5-30s


10r=5-30(-1/3)


10r=5+10

10r=15

r=15/10

r=3/2 <--------Answer # 2